5.Dawn Floral Creations, Inc., makes jewelry in the shape of flowers. Each piece is hand-made and takes an average of 1.5 hours to produce because of the intricate design and scrollwork. Dawn uses direct labor hours to allocate the overhead cost to production. Fixed overhead costs, including rent, depreciation, supervisory salaries, and other production expenses, are budgeted at $9,000 per month. Dawns budgeted output is 1,000 pieces, variable cost per unit is $25 and it expects to sell its jewelry for S55 a piece. During the month of February, Dawn produced 600 pieces of jewelry and actual fixed costs were $9,200. Calculate the operating income volume variance: A$8,600 U B$8,400 U C$18,000 F DS6,000 U
Answer #1
Budgeted variable cost per unit $ 25.00
Budgeted fixed cost per unit = 9000/1000 = $    9.00
Budgeted cost per unit $ 34.00
Sale price per unit $ 55.00
Budgeted net operating income per unit $ 21.00
Operating income volume variance = 21*(1000-600) = $ 8,400 U